# Preventing mass loss in the standard level set method: New insights from variational analyses

### v1

@@ 2.2 Interface tracking
It is well known that the level set field does not maintain the signed distance property after advection [1, 14]. Retaining the signed distance property is essential as the advected $\phi$ is used to prescribe various material properties for the one-fluid model and calculate geometric quantities like the interface curvature and surface normals. A reinitialization step is typically performed after the advection step to regain the signed distance property of $\phi$. The reinitialization step involves time-advancing the Hamilton-Jacobi equation to steady state
@@ 2.3 Precise reason for mass loss with the standard level set method
Leibniz rule for a moving domain: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}(t)}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}(t)}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \rm{d}\mathcal{V}+\int_{\mathcal{A}(t)}\mathit{f}\leavevmode\nobreak\ (\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{d}\mathcal{A},$ (7a)
Leibniz rule for a static domain: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \rm{d}\mathcal{V},$ (7b)
@@
RTT for a material volume: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}(t)}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}(t)}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \rm{d}\mathcal{V}+\int_{\mathcal{A}(t)}\mathit{f}\leavevmode\nobreak\ (\bm{\mathrm{u}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \rm{d}\mathcal{A}.$ (8)
@@
$\Delta M$ $=(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)(-\breve{\bm{\mathrm{u}}}\cdot\nabla\phi)\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \rm{d}t,$
$=(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)(-\breve{\bm{\mathrm{u}}}\cdot|\nabla\phi|\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \rm{d}t,$
$=-(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)\,\breve{u}_{\text{n}}\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \rm{d}t.$ (10)
@@
Proceeding analogously, one can conserve the mass of fluid-2 $M_{2}=\int_{\Omega}\rho_{2}(1-H(\phi))\leavevmode\nobreak\ \text{d}\Omega$ by imposing a constraint of the form $\mathcal{C}(\phi)\overset{\triangle}{=}\int_{\Omega}\rho_{2}(1-H(\phi))-M^{0}_{2}/V\leavevmode\nobreak\ \text{dV}=0$ with the help of Lagrange multiplier $\lambda_{2}\in L^{2}(\Omega)^{3}$. In this case the dynamical equation for the level set reads as
@@ 2.6 Conserving mass is the same as conserving volume for incompressible flows
$\frac{\rm{d}V_{1}}{\rm{d}t}$ $=\frac{\text{d}}{\text{d}t}\int_{\Omega}H(\phi)\leavevmode\nobreak\ \text{dV}=\int_{\Omega}\delta(\phi)\frac{\partial\phi}{\partial t}\leavevmode\nobreak\ \text{dV}.$ (23b)
@@
$\frac{\partial\phi}{\partial t}+\breve{\bm{\mathrm{u}}}\cdot\nabla\phi=\frac{\rm{d}\varepsilon}{\rm{d}t}.$ (28)
To relate the informal Eq. (28) and the formal Eq. (20), we first consider an intuitive argument and then present a mathematical derivation. Note that the “forcing” that induces a shift in the contour levels of $\phi$ is the RHS of Eq. (28). This term is taken to be uniform in the entire domain leading to the same correction to every contour level. The corresponding RHS of the formal Eq. (20) is $\beta\delta(\phi)$, where $\beta$ is uniform everywhere. However, the $\delta(\phi)$ term makes the “forcing” in the formal equation non-uniform in the domain. If one picks a contour level, say $\phi=0$, then the “forcing” does have the same value on that entire contour and this will lead to the same correction to $\phi$ on that contour. Similarly, each contour level has its own constant value of correction to $\phi$. This correction is maximum on $\phi=0$ contour and decreases away from it until it is zero outside the smeared interfacial region where $\delta(\phi)=0$. Since, in the formal method of Eq. (20) each contour level is corrected by a different amount, the distance function property is lost. The only way to preserve the distance function property of $\phi$ is to correct all contour levels by the same amount. This is the assumption made in the informal approach of Eq. (28). Thus, Eq. (28) would arise from Eq. (20) under the assumption that $\frac{\rm{d}\varepsilon}{\rm{d}t}=\beta\delta(\phi_{m})$, which means that the correction of an appropriately chosen $\phi_{m}$ contour in the formal approach is uniformly applied in the entire domain to obtain the informal approach.
We now proceed with the mathematical derivation by obtaining an expression for the rate of change of correction $\frac{\rm{d}\varepsilon}{\rm{d}t}$. This can be obtained by differentiating Eq. (26) with respect to time
@@
$\delta(\phi)=\frac{1}{2\Delta}+\frac{1}{2\Delta}\cos\left(\frac{\pi\phi}{\Delta}\right)\quad\hookrightarrow\int_{-\Delta}^{\Delta}\delta(x)\,\rm{d}x=1\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \int_{-\Delta}^{\Delta}\delta^{2}(x)\,\rm{d}x=\frac{3}{4\Delta},$ (31)
@@ 2.8 Comparison with other mass-preserving level set methods
Here, $L_{*}(\phi)=[(\rho_{1}-2\rho_{2})+2(\rho_{2}-\rho_{1})H(\phi)]\delta(\phi)$. Note that Eq. (41) differs from Eq. (40). Each fluid has its own governing equation. As part of their numerical experiments, Wen et al. (arbitrarily) chose the level set equation to preserve fluid-1’s mass. The only way to conserve both fluids’ masses simultaneously with this approach is to have $L_{1}(\phi)=L_{*}(\phi)$. This condition simplifies to yield $4(\rho_{1}-\rho_{2})H(\phi)=\rho_{1}-3\rho_{2}$. Considering that the RHS is constant, and the LHS varies spatially, this is a contradiction. There is also an unphysical condition $\rho_{1}=-\rho_{2}$ at the interface where $\phi=0$ and $H(\phi=0)=1/2$.
@@ 2.9.1 Conserving mass/volume for three phase flows - Non-immersed formulation
$\hat{\delta}^{\text{total}}\mathcal{L}(\Lambda_{1},\phi)=\hat{\delta}\mathcal{L}(\Lambda_{1},\phi)+\Lambda_{1}\int_{\partial\Omega^{\text{S}}(t)}\left(H(\phi)-\frac{V^{0}_{1}}{V}\right)(\bm{\mathrm{v}}_{\text{s}}\rm{d}t\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dS},$
@@
which is iterated till the relative error ($\Delta=f(\varepsilon_{k})/V^{0}_{1}$) drops to machine accuracy. To show the similarity between the exact and approximate Lagrange multiplier approaches for three phase flows, we need an expression for level set correction per unit time $\frac{\rm{d}\varepsilon}{\rm{d}t}$ instead of the total one: $\varepsilon$. This can obtained by differentiating $f(\varepsilon)$, which is a non conserved quantity defined over a moving domain, with respect to time using the Leibniz integral rule
@@ 2.9.2 Conserving mass/volume for three phase flows - Immersed formulation
The value of the Lagrange multiplier $\Lambda_{3}$ in Eq. (58b) is obtained by substituting $\frac{\partial\phi}{\partial t}=-\breve{\bm{\mathrm{u}}}\cdot\nabla\phi-\Lambda_{3}\delta(\phi)H(\zeta)$ from Eq. (58a) and $\partial\zeta/\partial t=-\bm{\mathrm{u}}\cdot\nabla\zeta$ from Eq. (45) into Eq. (56a) and setting $\frac{\rm{d}V_{1}}{\rm{d}t}=0$. If the body delta function $\delta(\zeta)$ is sharp, then the expression for $\Lambda_{3}$ in Eq. (58b) can be further simplified to Eq. (58c) by using the relation $\int_{\Omega}H(\phi)\delta(\zeta)(\bm{\mathrm{u}}\cdot\nabla\zeta)=\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\nabla\zeta)\leavevmode\nobreak\ \text{dS}=\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}}_{\rm s})\leavevmode\nobreak\ \text{dS}=-\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dS}$. Here, $\bm{\mathrm{n}}_{\rm s}=-\bm{\mathrm{n}}$ is the outward unit normal to the solid surface, which can be obtained from the signed distance function $\zeta$ as $\bm{\mathrm{n}}_{\rm s}=\nabla\zeta/|\nabla\zeta|=\nabla\zeta$.
@@ 2.9.4 Differential treatment of the Brinkman penalty force
    the system of equations become stiff to solve; and
@@
    a no-slip condition in the tangential direction implies a stress singularity at the material triple points.
@@ 3.3.2 Multiphase incompressible Navier-Stokes solution
in which, $\bm{\mathrm{C}}^{n+1,k}$ is the discrete version of the convective term $\nabla\cdot\mbox{}(\bm{\mathrm{m_{\rho}}}\otimes\bm{\mathrm{u}})$. A consistent mass/momentum transport scheme is used to compute the density $\breve{\bm{\mathrm{\rho}}}^{n+1,k+1}$ and mass flux $\bm{\mathrm{m_{\rho}}}\equiv\rho\bm{\mathrm{u}}$ approximations, ensuring numerical stability even under high density contrasts. Our prior works [21, 17] provide a detailed description of the consistent mass/momentum transport scheme. The viscous strain rate in Eq. (71) is calculated using the Crank-Nicolson approximation: $\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n+\frac{1}{2},k+1}=\frac{1}{2}\left[\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n+1,k+1}+\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n}\right]$, in which $\left(\bm{\mathrm{L}}_{\mu}\right)^{n+1}=\nabla_{h}\cdot\left[\mu^{n+1}\left(\nabla\bm{\mathrm{u}}+\nabla\bm{\mathrm{u}}^{T}\right)^{n+1}\right]$. The newest approximation to viscosity $\mu^{n+1,k+1}$ is obtained using the two-stage process described by Eqs. (65) and (66).

### v2

@@ 2.2 Interface tracking
It is well known that the level set field does not maintain the signed distance property after advection [1, 14]. Although this fact is widely known, its proof is not readily available in the literature (to our knowledge). The following derivation shows that $\phi$’s signed distance property is lost at the continuous level under advection. This is not purely a numerical artifact.
By taking gradient of Eq. (3) we get
$\frac{\partial\nabla\phi}{\partial t}+\nabla\left(\bm{\mathrm{u}}\cdot\nabla\phi\right)=0,$
$\hookrightarrow$ $\frac{\partial\nabla\phi}{\partial t}+\nabla\bm{\mathrm{u}}\cdot\nabla\phi+\bm{\mathrm{u}}\cdot\nabla\left(\nabla\phi\right)=0.$ (6)
Defining ${\bf{\Psi}}\equiv\nabla\phi$ in Eq. (6), and taking a dot product with ${\bf{\Psi}}$ on both sides, we obtain
${\bf{\Psi}}\cdot\left[\frac{\partial{\bf{\Psi}}}{\partial t}+\nabla\bm{\mathrm{u}}\cdot{\bf{\Psi}}+\bm{\mathrm{u}}\cdot\nabla{\bf{\Psi}}\right]=0,$
$\hookrightarrow$ ${\bf{\Psi}}\cdot\frac{\partial{\bf{\Psi}}}{\partial t}+{\bf{\Psi}}\cdot\nabla{\bf{\Psi}}\cdot\bm{\mathrm{u}}=-{\bf{\Psi}}\cdot\nabla\bm{\mathrm{u}}\cdot{\bf{\Psi}}.$ (7)
Using the relations ${\bf{\Psi}}\cdot\partial{\bf{\Psi}}/\partial t=\frac{1}{2}\frac{\partial({\bf{\Psi}}\cdot{\bf{\Psi}})}{\partial t}=|\nabla\phi|\frac{\partial|\nabla\phi|}{\partial t}$ and ${\bf{\Psi}}\cdot\nabla{\bf{\Psi}}\cdot\bm{\mathrm{u}}=\frac{\bm{\mathrm{u}}\cdot\nabla({\bf{\Psi}}\cdot{\bf{\Psi}})}{2}=|\nabla\phi|\bm{\mathrm{u}}\cdot\nabla(|\nabla\phi|)$, and simplifying further yields a transport equation for $|\nabla\phi|$
$\frac{\partial|\nabla\phi|}{\partial t}+\bm{\mathrm{u}}\cdot\nabla\left(|\nabla\phi|\right)=\frac{-\nabla\phi\cdot\nabla\bm{\mathrm{u}}\cdot\nabla\phi}{|\nabla\phi|}.$ (8)
The right hand side of Eq. (8) has a non-zero value if the symmetric part of the velocity gradient tensor, $\bm{\mathrm{S}}=\frac{1}{2}(\nabla\bm{\mathrm{u}}+\nabla\bm{\mathrm{u}}^{\intercal})$, is not zero. When the velocity field is a rigid body motion (translation and rotation), the right hand side of Eq. (8) will be zero, and $\phi$ will preserve its signed distance property under advection. However, in general, an advective transport of $\phi$ disrupts its signed distance property.
Retaining the signed distance property is numerically advantageous as the advected $\phi$ is used to prescribe various material properties for the one-fluid model and to calculate geometric quantities like the interface curvature and surface normals. To regain the signed distance property of $\phi$, a reinitialization step is typically performed after the advection step. The reinitialization step involves time-advancing the Hamilton-Jacobi equation to steady state
@@ 2.3 Precise reason for mass loss with the standard level set method
Leibniz rule for a moving domain: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}(t)}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}(t)}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \text{d}\mathcal{V}+\int_{\mathcal{A}(t)}\mathit{f}\leavevmode\nobreak\ (\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{d}\mathcal{A},$ (10a)
Leibniz rule for a static domain: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \text{d}\mathcal{V},$ (10b)
@@
RTT for a material volume: $\frac{\text{d}}{\text{d}t}\int_{\mathcal{V}(t)}f\leavevmode\nobreak\ \text{d}\mathcal{V}=\int_{\mathcal{V}(t)}\frac{\partial f}{\partial t}\leavevmode\nobreak\ \text{d}\mathcal{V}+\int_{\mathcal{A}(t)}\mathit{f}\leavevmode\nobreak\ (\bm{\mathrm{u}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{d}\mathcal{A}.$ (11)
@@
$\Delta M$ $=(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)(-\breve{\bm{\mathrm{u}}}\cdot\nabla\phi)\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \text{d}t,$
$=(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)(-\breve{\bm{\mathrm{u}}}\cdot|\nabla\phi|\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \text{d}t,$
$=-(\rho_{1}-\rho_{2})\int_{\Delta t}\int_{\Omega}\delta(\phi)\,\breve{u}_{\text{n}}\leavevmode\nobreak\ \text{dV}\leavevmode\nobreak\ \text{d}t.$ (13)
@@
Proceeding analogously, one can conserve the mass of fluid-2 $M_{2}=\int_{\Omega}\rho_{2}(1-H(\phi))\leavevmode\nobreak\ \text{dV}$ by imposing a constraint of the form $\mathcal{C}(\phi)\overset{\triangle}{=}\int_{\Omega}\rho_{2}(1-H(\phi))-M^{0}_{2}/V\leavevmode\nobreak\ \text{dV}=0$ with the help of Lagrange multiplier $\lambda_{2}\in L^{2}(\Omega)^{3}$. In this case the dynamical equation for the level set reads as
@@ 2.6 Conserving mass is the same as conserving volume for incompressible flows
$\frac{\text{d}V_{1}}{\text{d}t}$ $=\frac{\text{d}}{\text{d}t}\int_{\Omega}H(\phi)\leavevmode\nobreak\ \text{dV}=\int_{\Omega}\delta(\phi)\frac{\partial\phi}{\partial t}\leavevmode\nobreak\ \text{dV}.$ (26b)
@@
$\frac{\partial\phi}{\partial t}+\breve{\bm{\mathrm{u}}}\cdot\nabla\phi=\frac{\text{d}\varepsilon}{\text{d}t}.$ (31)
To relate the informal Eq. (31) and the formal Eq. (23), we first consider an intuitive argument and then present a mathematical derivation. Note that the “forcing” that induces a shift in the contour levels of $\phi$ is the RHS of Eq. (31). This term is taken to be uniform in the entire domain leading to the same correction to every contour level. The corresponding RHS of the formal Eq. (23) is $\beta\delta(\phi)$, where $\beta$ is uniform everywhere. However, the $\delta(\phi)$ term makes the “forcing” in the formal equation non-uniform in the domain. If one picks a contour level, say $\phi=0$, then the “forcing” does have the same value on that entire contour and this will lead to the same correction to $\phi$ on that contour. Similarly, each contour level has its own constant value of correction to $\phi$. This correction is maximum on $\phi=0$ contour and decreases away from it until it is zero outside the smeared interfacial region where $\delta(\phi)=0$. Since, in the formal method of Eq. (23) each contour level is corrected by a different amount, the distance function property is lost. The only way to preserve the distance function property of $\phi$ is to correct all contour levels by the same amount. This is the assumption made in the informal approach of Eq. (31). Thus, Eq. (31) would arise from Eq. (23) under the assumption that $\frac{\text{d}\varepsilon}{\text{d}t}=\beta\delta(\phi_{m})$, which means that the correction of an appropriately chosen $\phi_{m}$ contour in the formal approach is uniformly applied in the entire domain to obtain the informal approach.
We now proceed with the mathematical derivation by obtaining an expression for the rate of change of correction $\frac{\text{d}\varepsilon}{\text{d}t}$. This can be obtained by differentiating Eq. (29) with respect to time
@@
$\delta(\phi)=\frac{1}{2\Delta}+\frac{1}{2\Delta}\cos\left(\frac{\pi\phi}{\Delta}\right)\quad\hookrightarrow\int_{-\Delta}^{\Delta}\delta(x)\,\text{d}x=1\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \int_{-\Delta}^{\Delta}\delta^{2}(x)\,\text{d}x=\frac{3}{4\Delta},$ (34)
@@ 2.8 Comparison with other mass-preserving level set methods
Here, $L_{*}(\phi)=[(\rho_{1}-2\rho_{2})+2(\rho_{2}-\rho_{1})H(\phi)]\delta(\phi)$. Note that Eq. (44) differs from Eq. (43). Each fluid has its own governing equation. As part of their numerical experiments, Wen et al. (arbitrarily) chose the level set equation to preserve fluid-1’s mass. The only way to conserve both fluids’ masses simultaneously with this approach is to have $L(\phi)=L_{*}(\phi)$. This condition simplifies to yield $4(\rho_{1}-\rho_{2})H(\phi)=\rho_{1}-3\rho_{2}$. Considering that the RHS is constant, and the LHS varies spatially, this is a contradiction. There is also an unphysical condition $\rho_{1}=-\rho_{2}$ at the interface where $\phi=0$ and $H(\phi=0)=1/2$.
@@ 2.9.1 Conserving mass/volume for three phase flows - Non-immersed formulation
$\hat{\delta}^{\text{total}}\mathcal{L}(\Lambda_{1},\phi)=\hat{\delta}\mathcal{L}(\Lambda_{1},\phi)+\Lambda_{1}\int_{\partial\Omega^{\text{S}}(t)}\left(H(\phi)-\frac{V^{0}_{1}}{V}\right)(\bm{\mathrm{v}}_{\text{s}}\text{d}t\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dS},$
@@
which is iterated till the relative error ($\Delta=f(\varepsilon_{k})/V^{0}_{1}$) drops to machine accuracy. To show the similarity between the exact and approximate Lagrange multiplier approaches for three phase flows, we need an expression for level set correction per unit time $\frac{\text{d}\varepsilon}{\text{d}t}$ instead of the total one: $\varepsilon$. This can obtained by differentiating $f(\varepsilon)$, which is a non conserved quantity defined over a moving domain, with respect to time using the Leibniz integral rule
@@ 2.9.2 Conserving mass/volume for three phase flows - Immersed formulation
The value of the Lagrange multiplier $\Lambda_{3}$ in Eq. (61b) is obtained by substituting $\frac{\partial\phi}{\partial t}=-\breve{\bm{\mathrm{u}}}\cdot\nabla\phi-\Lambda_{3}\delta(\phi)H(\zeta)$ from Eq. (61a) and $\partial\zeta/\partial t=-\bm{\mathrm{u}}\cdot\nabla\zeta$ from Eq. (48) into Eq. (59a) and setting $\frac{\text{d}V_{1}}{\text{d}t}=0$. If the body delta function $\delta(\zeta)$ is sharp, then the expression for $\Lambda_{3}$ in Eq. (61b) can be further simplified to Eq. (61c) by using the relation $\int_{\Omega}H(\phi)\delta(\zeta)(\bm{\mathrm{u}}\cdot\nabla\zeta)=\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\nabla\zeta)\leavevmode\nobreak\ \text{dS}=\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}}_{\rm s})\leavevmode\nobreak\ \text{dS}=-\int_{\partial\Omega^{\text{S}}(t)}H(\phi)(\bm{\mathrm{u}}_{\text{s}}\cdot\bm{\mathrm{n}})\leavevmode\nobreak\ \text{dS}$. Here, $\bm{\mathrm{n}}_{\rm s}=-\bm{\mathrm{n}}$ is the outward unit normal to the solid surface, which can be obtained from the signed distance function $\zeta$ as $\bm{\mathrm{n}}_{\rm s}=\nabla\zeta/|\nabla\zeta|=\nabla\zeta$.
@@ 2.9.4 Differential treatment of the Brinkman penalty force
    the system of equations becomes stiff to solve; and
@@
    a no-slip condition in the tangential direction implies stress singularity at the material triple points.
@@ 3.3.2 Multiphase incompressible Navier-Stokes solution
in which, $\bm{\mathrm{C}}^{n+1,k}$ is the discrete version of the convective term $\nabla\cdot\mbox{}(\bm{\mathrm{m_{\rho}}}\otimes\bm{\mathrm{u}})$. A consistent mass/momentum transport scheme is used to compute the density $\breve{\bm{\mathrm{\rho}}}^{n+1,k+1}$ and mass flux $\bm{\mathrm{m_{\rho}}}\equiv\rho\bm{\mathrm{u}}$ approximations, ensuring numerical stability even under high density contrasts. Our prior works [21, 17] provide a detailed description of the consistent mass/momentum transport scheme. The viscous strain rate in Eq. (74) is calculated using the Crank-Nicolson approximation: $\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n+\frac{1}{2},k+1}=\frac{1}{2}\left[\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n+1,k+1}+\left(\bm{\mathrm{L}}_{\mu}\bm{\mathrm{u}}\right)^{n}\right]$, in which $\left(\bm{\mathrm{L}}_{\mu}\right)^{n+1}=\nabla_{h}\cdot\left[\mu^{n+1}\left(\nabla\bm{\mathrm{u}}+\nabla\bm{\mathrm{u}}^{T}\right)^{n+1}\right]$. The newest approximation to viscosity $\mu^{n+1,k+1}$ is obtained using the two-stage process described by Eqs. (68) and (69). To avoid generating spurious currents due to large density variations near the fluid-solid interface [22], the gravitational body force term $\bm{\mathrm{\wp}}\bm{\mathrm{g}}=\bm{\mathrm{\rho}}^{\text{flow}}\bm{\mathrm{g}}$ is calculated using the flow density field; see also Eq. (68).